POTENTIAL ENERGY SURFACES AND BRANCHING RATIO OF THE DISSOCIATIVE RECOMBINA-TION REACTION HCNH^{+} + e^{−} : AN AB INITIO MOLECULAR ORBITAL STUDY

\underline{TSUNEO HIRANO}, and YOKO SHIBA, Department of Chemistry, Faculty of Science, Ochanomizu University, 2-1-1 Otsuka, Bunkyo-ku, Tokyo 112, Japan; UMPEI NAGASHIMA, Department of Information Sciences, Faculty of Science, Ochanomizu University, 2-1-1 Otsuka, Bunkyo-ku, Tokyo 112, Japan; KEISAKU ISHII, Department of Applied Chemistry, Faculty of Engineering, University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113, Japan.

Two  dimensional  potential  energy  surfaces  for  the  dissociative  recombination  reaction,

HCNH^{+} + e^{−} → (HCNH)^{∗} → HCN/HNC + H, have been calculated by the CASSCF-MRSDCI/cc-pVQZ ab initio molecular orbital method. Near the HCNH^{+} ground state surface, which also represents the upper limit of the neutral HCNH Rydberg state surface, only two dissociative valence states of the HCNH are located: One is the ^{2}Σ^{+} state resulting in the H-C bond scission yielding HNC and the other the ^{2}Σ^{+} state resulting in the H-N bond scission yielding HCN. The seam of intersection of these two surfaces almost bisects the zero-point vibrational wave function of the Rydberg-limit state of the neutral HCNH species, indicating that, while descending Rydberg ladder, transition to each dissociative surface occurs with almost equal probability. Thus, the branching ratio [HNC]/[HCN] is predicted to be nearly one or slightly more, explaining the thermochemically unrealistic interstellar [HNC]/[HCN] ratio of 4.4 observed toward L134N.